0.5x+y=9
1.6x+0.2y=13
Can you show me the steps to solve this using elimination?

Answer by Alan3354(69443)

0.5x+y=9
1.6x+0.2y=13
----------
Multiply the 1st eqn by 2 and the 2nd by 10. --> no fractions
-----
1x + 2y = 18
16x + 2y = 130
---------------- Subtract
-15x = -112
x = 112/15
-----
x + 2y = 18
112/15 + 2y = 18
y = 79/15